Daily sales record for a computer store show that it will sell 1, 2, or 3 computers with the probability as listed:

Numbers of Sales 1 2 3
Probability 0.3 .06 .01

A) What is the expected value and variance of daily sales?

B) In the table below, compute the probability of X, the number of sales in a two-day period, assuming that sales are independent from day to day.

Number of Sales (X) 2 3 4 5 6
Probability
Answer #1

A)
Expected value = 1*0.3 + 2*0.06 + 3*0.01 = 0.4500

B)
for X = 2, probability = 0.3*0.3 = 0.09
for X = 3, probability = 2*(0.3*0.06) = 0.0360
for X = 4, probability = 2*0.3*0.01 + 0.06*0.06 = 0.0096
for X = 5, proability = 2*0.06*0.01 = 0.0012
for X = 6, probability = 0.01*0.01 = 0.0001